The locus of the middle points of the focal chords of the parabola ,y^2=4x isAns=y^2=2(x-1)Plz give the full solution 😒😒

Grade:

1 Answers

Arun
6 years ago
Β 

Let the parabola we consider and draw chords be y2Β = 4ax.

The Vertex is O(0.0), which is one end of the chord. Let the other end be a varaible point P given by (at2,2at).

Β 

Let M(p,q) be the midpoint of the chord OP.

Midpoint of OP is (at2/2,at).

So, p = at2/2 and q = at. Now we have to eliminate "t" and get the relation between p and q to get the locus.

So t = q/a. Substitute this in the equation of p, and we will get

p = a/2*(q/a)2

So we have q2Β = 2ap.

Β 

Which is a parabola of the form y2Β = 2ax. And that proves the result.